ReviewI ConstantsI Periodic Table Use the inverse square law for light to answer each of the following questions Part C Suppose a star has a luminosity of 6.0x1028 watts and an apparent brightness of 2.0x10-12 watt/m2. How far away is i? Give your answer in both kilometers and light-years Express your answer using two significant figures km Submit Part D Express your answer using two significant figures d- light - years Submit Request Answer
Answer #1

We know that the brightness is given by
B = rac{L}{4pi D^{2}}
Where B is brightness
L is luminosity
D is distance
Now on putting the values
2*10-12 = 6#1 4Td2
d = 4.89*1018 m
We know that
1 km = 1000 m , therefore
d = 4.89*1015 km
Now we know that
1 light year = 9.461*1012 km
d = 4.89*1015*(1/9.461*1012 ) = 516.44 light years